The spot rate of the British pound to the dollar is $1.15. The 180-day forward rate is $1.17. Thus, the approximate annualized forward premium is _________.
Answer #1

The forward Premium = (Forward Rate/Spot Rate-1)*365/No of days =(1.17/1.15 -1)*365/180 = 3.53%

If 360 days convention is used then forward premium =(Forward Rate/Spot Rate-1)*360/No of days =(1.17/1.15 -1)*360/180 = 3.48%